# Exploring Video-Based Driver Activity Recognition under Noisy Labels

### v1

@@ II-B Driver Activity Recognition
Deep learning has shown promising performances on various of areas, e.g., natural language processing [24, 25], video understanding [26, 27], and human activity recognition [28, 29, 30]. Driver activity recognition is a subdirection under human activity recognition, which also moves from the approaches relies on handcrafted features into end-to-end learning pipelines [31, 32, 33]. Traditional driver behavior recognition relies on handcrafted features with classifiers like SVMs [34] or random forests [35]. With the rise of CNNs [36], deep learning pipelines dominate driver activity analysis, employing top CNNs [37], spatiotemporal models like I3D [38]. Recent works explore advanced techniques to achieve driver activity recognition. Zhao et al.[39] adopt spatial attention mechanisms, while Peng et al. [4] illustrate the strength of transformers when dealing with the driver activity recognition task. In this work, we build the first benchmark to achieve noisy label learning in the driver activity recognition field. Unlike conventional noisy label learning problems that primarily deal with classification noise in static images, our approach is tailored to handle noisy annotations in video-based tasks, incorporating spatiotemporal contrastive learning and clustering-based refinement to enhance feature learning under label noise. Additionally, we incorporate a class-balancing strategy to mitigate the long-tailed distribution of driver activities, a critical aspect of this domain. These design choices explicitly address the complexities of driver behavior recognition.
@@ III-A1 Spatiotemporal Contrastive Learning
Additionally, we utilize a linear decoder $\mathbf{W}_{\mathrm{d}}$ to reconstruct the high-dimensional embeddings from the low-dimensional embeddings. An L2 reconstruction loss is introduced to preserve the mutual information between the latent spaces, defined as:
@@ III-A2 Label Noise Agnostic Clustering Module
Analogous to the instance discrimination method, a softmax-formulated constraint is applied to encourage the correlation between a feature vector and itself to converge to $1$, while the correlation between a feature vector and other vectors converges to ${-}1$ or 00 [44]. The constraint is defined as:
@@ 16       $\theta=$SGD $(\mathcal{L},\theta)$
20      $\{\mathbf{c}_{i}\}_{i=1}^{K}=$Clustering-Algorithm ($(\{\mathbf{v}_{i}\}_{i=1}^{n})$
21      $\{\mathbf{p}_{i}^{t}\}_{i=1}^{n}=$softmax ($(\{f_{\theta}(\mathbf{x}_{i})\}_{i=1}^{n})$
@@ III-B3 Clean Samples Selection
where $X^{t-1}(c)$ indicates the clean subset containing the filtered samples with $y_{i}{=}c$ from the epoch $t{-}1$. $D(c)$ denotes the set of training samples with the observed label $c$. $\epsilon$ is a small constant introduced to prevent division by zero and ensure numerical stability.
@@ III-B4 Soft pseudo-labels generation for noisy samples
$\tilde{q}_{i}^{t}=\begin{cases}\hat{p}_{i,k}^{t}&\text{ if }k\in\{\mathbf{y}_{i},\mathbf{y}_{i}^{*}\}\\ 0&\text{ otherwise. }\end{cases}$ (14)
@@ IV-A Evaluation Results on Drive&Act Dataset
Implementation details. We employ either a transformer-based video model MViTv2 [3] or a convolution-based video model X3D [2] as the video backbone for representation encoding, following the official code of PySlowFast [^1] https://github.com/facebookresearch/SlowFast. We employ the AdamW optimizer [46] with an initial learning rate of $41e{-}$ 4 and adopt a cosine annealing learning rate scheduler. For our method, we build an additional linear projection layer with the output dimensionality $d{=}48$. We sample two clips out of $16$ frames with a sampling rate of $4$ from each input raw video and apply a random combination of spatial augmentations, including random $224^{2}$ crop, horizontal flip, color distortion, and Gaussian blur to all the sampled clips. We warm up the model for $5$ epochs and apply an agglomerative clustering algorithm [47] to assign the learned representations into clusters. For further details in hyperparameter tuning, we set $\tau_{1}{=}0.3,\tau_{2}{=}0.5,\tau_{3}{=}2.0$, $\alpha{=}8.0$. We tune $\beta{=}0.5$ for the co-refinement within a cluster and set $\epsilon{=}0.05$. The weight of L2 loss on the pseudo-labeled subset is initially defined as $\gamma_{u}{=}100$ for MViTv2, while set as $\gamma_{u}{=}50$ for X3D. All the training is conducted on an NVIDIA Quadro RTX 8000 graphic card with 50GB RAM with a batch size of $4$. All the hyperparameter selections are conducted through griding search.
@@ V Conclusion
We propose a robust learning framework for video-based driver activity recognition under noisy labels, combining clustering-friendly representation learning, hyperparameter-free sample selection, and adaptive class balancing. Our method outperforms existing noisy label approaches on the Drive&Act benchmark across fine-grained, coarse, and structured recognition tasks. Ablation studies validate the effectiveness of each component in enhancing robustness and generalization. This work offers an efficient and noise-resilient solution for real-world driver behavior analysis, with potential for broader applications in intelligent transportation systems. Future research may extend the framework to multi-modal or cross-domain scenarios.

### v2

@@ II-B Driver Activity Recognition
Deep learning has shown promising performances on various of areas, e.g., natural language processing [24, 25], video understanding [26, 27], and human activity recognition [28, 29, 30]. Driver activity recognition is a subdirection under human activity recognition, which also moves from the approaches relying on handcrafted features into end-to-end learning pipelines [31, 32, 33]. Traditional driver behavior recognition relies on handcrafted features with classifiers like SVMs [34] or random forests [35]. With the rise of CNNs [36], deep learning pipelines dominate driver activity analysis, employing top CNNs [37], spatiotemporal models like I3D [38]. Recent works explore advanced techniques to achieve driver activity recognition. Zhao et al.[39] adopt spatial attention mechanisms, while Peng et al. [4] illustrate the strength of transformers when dealing with the driver activity recognition task. In this work, we build the first benchmark to achieve noisy label learning in the driver activity recognition field. Unlike conventional noisy label learning problems that primarily deal with classification noise in static images, our approach is tailored to handle noisy annotations in video-based tasks, incorporating spatiotemporal contrastive learning and clustering-based refinement to enhance feature learning under label noise. Additionally, we incorporate a class-balancing strategy to mitigate the long-tailed distribution of driver activities, a critical aspect of this domain. These design choices explicitly address the complexities of driver behavior recognition.
@@ III-A1 Spatiotemporal Contrastive Learning
Additionally, we utilize a linear decoder $\mathbf{W}_{\mathrm{d}}$ to reconstruct the high-dimensional embedding $\mathbf{u}_{i}=f_{\theta}(\mathbf{x}_{i})$ from the low-dimensional representation $\mathbf{v}_{i}$, where $f_{\theta}$ denotes the video encoder parameterized by $\theta$. The decoder aims to preserve the mutual information between the latent spaces. The L2 reconstruction loss is defined as:
@@ III-A2 Label Noise Agnostic Clustering Module
Analogous to the instance discrimination method, a softmax-formulated constraint is applied to encourage the correlation between a feature vector and itself to converge to $1$, while the correlation between a feature vector and other vectors converges to 1{-}1 or 0 [44]. The constraint is defined as:
@@ 16 $\theta=$SGD $(\mathcal{L},\theta)$
20 $\{\mathbf{c}_{i}\}_{i=1}^{K}=$Clustering-Algorithm $(\{\mathbf{v}_{i}\}_{i=1}^{n})$
21 $\{\mathbf{p}_{i}^{t}\}_{i=1}^{n}=$softmax $(\{f_{\theta}(\mathbf{x}_{i})\}_{i=1}^{n})$
@@ III-B3 Clean Samples Selection
where $X^{t-1}(c)$ indicates the clean subset containing the filtered samples with $y_{i}{=}c$ from the epoch t1t{-}1. $D(c)$ denotes the set of training samples with the observed label $c$. $\epsilon$ is a small constant introduced to prevent division by zero and ensure numerical stability.
@@ III-B4 Soft pseudo-labels generation for noisy samples
[…]
@@ IV-A Evaluation Results on Drive&Act Dataset
Implementation details. We employ either a transformer-based video model MViTv2 [3] or a convolution-based video model X3D [2] as the video backbone for representation encoding, following the official code of PySlowFast [^1] https://github.com/facebookresearch/SlowFast. We employ the AdamW optimizer [46] with an initial learning rate of 1e41e{-}4 and adopt a cosine annealing learning rate scheduler. For our method, we build an additional linear projection layer with the output dimensionality d=48d{=}48. We sample two clips out of $16$ frames with a sampling rate of $4$ from each input raw video and apply a random combination of spatial augmentations, including random $224^{2}$ crop, horizontal flip, color distortion, and Gaussian blur to all the sampled clips. We warm up the model for $5$ epochs and apply an agglomerative clustering algorithm [47] to assign the learned representations into clusters. For further details in hyperparameter tuning, we set $\tau_{1}{=}0.3,\tau_{2}{=}0.5,\tau_{3}{=}2.0$, $\alpha{=}8.0$. We tune $\beta{=}0.5$ for the co-refinement within a cluster and set $\epsilon{=}0.05$. The weight of L2 loss on the pseudo-labeled subset is initially defined as $\gamma_{u}{=}100$ for MViTv2, while set as $\gamma_{u}{=}50$ for X3D. All the training is conducted on an NVIDIA Quadro RTX 8000 graphic card with 50GB RAM with a batch size of $4$. All the hyperparameter selections are conducted through griding search.
@@ V Conclusion
We propose a robust framework for video-based driver activity recognition under noisy labels, integrating clustering-friendly representation learning, hyperparameter-free sample selection, and adaptive class balancing. Our method outperforms existing approaches on the Drive&Act benchmark across various recognition tasks. Ablation studies confirm the effectiveness of each component. This efficient, noise-resilient solution supports real-world driver behavior analysis and can be extended to multi-modal or cross-domain applications.